# ReflecSched: Solving Dynamic Flexible Job-Shop Scheduling via LLM-Powered Hierarchical Reflection

### v1

@@ Abstract
Dynamic Flexible Job-Shop Scheduling (DFJSP) is an NP-hard problem challenged by real-time event adaptation and complex machine routing. While traditional dispatching rules are efficient but rigid, deep learning approaches are opaque and require intricate feature engineering. Large Language Models (LLMs) promise adaptive reasoning without this engineering overhead, yet we find their direct application is suboptimal. Baseline LLMs suffer from three key pitfalls: the long-context paradox, where crucial data is underutilized; an underutilization of expert heuristics; and myopic decision-making. To address this, we propose ReflecSched, a framework that empowers the LLM beyond a direct scheduler by equipping it with a strategic analysis capability. ReflecSched tasks the LLM to analyze heuristic-driven simulations across multiple planning horizons and distill them into a concise, natural-language summary termed “Strategic Experience”. This summary is then integrated into the prompt of a final decision-making module, guiding it to produce non-myopic actions. Experiments show that ReflecSched not only statistically significantly outperforms direct LLM baselines, securing a 71.35% Win Rate and a 2.755% Relative Percentage Deviation reduction, but also surpasses the performance of all individual heuristics evaluated, all while demonstrably mitigating the three identified pitfalls. Additionally, ReflecSched performs on par with the best heuristic tailored to each instance across all problem cases.
@@ 1 Introduction
Dynamic Flexible Job-Shop Scheduling (DFJSP) is a long-standing challenge in operations research and a key component of modern smart manufacturing (Li et al. 2025). As an NP-hard problem, it involves the complex task of continuously allocating operations to machines in a real-time environment characterized by stochastic events such as new job arrivals or machine breakdowns. The ability to efficiently schedule these operations critically impacts the agility, resilience, and profitability of production systems. For decades, the field has been dominated by methods ranging from classical heuristics to meta-heuristics (Cao et al. 2023). While foundational, these approaches often rely on handcrafted rules that can struggle to generalize across diverse and unforeseen dynamic scenarios (Ferreira, Figueira, and Amorim 2022).
In recent years, deep learning (DL), particularly reinforcement learning (RL), has emerged as a promising paradigm for this problem (Xu et al. 2025). By training agents on simulated data, DL-based methods can learn sophisticated, state-aware scheduling policies. However, this capability comes at a significant cost. A key limitation lies in the intricate process of state and action space engineering: translating the rich state of a dynamic workshop into a fixed-size numerical vector that a neural network can process is a complex task requiring dual expertise in scheduling and machine learning (Zhang et al. 2020). Furthermore, the resulting models are often “black boxes” with opaque decision-making logic that is difficult for engineers to interpret or trust in high-stakes environments (Ngwu, Liu, and Wu 2025).
The advent of Large Language Models (LLMs) offers a different approach, potentially circumventing these challenges by leveraging natural language as a more accessible interface. An LLM-based scheduler could, in principle, reason directly over a textual description of the factory state, reducing the need for complex numerical encoding (Abgaryan, Harutyunyan, and Cazenave 2024). This suggests the possibility of developing and adapting scheduling strategies with greater flexibility.
@@
  1. 1.
    We present a systematic investigation into the application of modern LLMs to the full scope of the DFJSP, incorporating a rich set of dynamic events.
  2. 2.
    We provide an empirical analysis that identifies and validates three key failure modes of baseline LLM-based schedulers, offering insights for the broader field of LLM-based planning.
  3. 3.
    We propose ReflecSched, a framework that addresses these pitfalls through a combination of hierarchical simulation and LLM-driven experience synthesis.
@@ 2 Related Work
Heuristic-based methods are a foundational approach for the Dynamic Job-Shop Scheduling Problem (DJSP) due to their computational tractability. Research has advanced from simple dispatching rules to complex metaheuristics designed to handle diverse operational constraints. For instance, some works have applied GRASP algorithms to manage a wide array of dynamic events (Baykasoğlu, Madenoğlu, and Hamzadayı 2020), while others have used hybrid Particle Swarm Optimization for joint production and transportation scheduling (Ren et al. 2022). Further research has focused on specific uncertainties, such as variable processing times, using techniques like Artificial Bee Colony algorithms (Shahgholi Zadeh, Katebi, and Doniavi 2019). A key limitation of these metaheuristics, however, is their reliance on a pre-defined search logic that is inherently less adaptive and struggles to generalize to new, unseen problem scenarios.
Deep Reinforcement Learning Approaches.
@@
To address the generalization limitations of heuristics, Deep Reinforcement Learning (DRL) has emerged as an alternative that learns adaptive policies, typically by formulating the DJSP as a Markov Decision Process. Architectural innovations have been central to this area. For example, Pointer Networks have been utilized for end-to-end policy learning that avoids reliance on predefined rules (Yang et al. 2024). To enhance scalability, a multi-agent perspective has been adopted through decentralized frameworks (Liu, Piplani, and Toro 2023). Other approaches decompose the decision-making process using hierarchical structures (Luo, Zhang, and Fan 2021). Nevertheless, DRL presents its own significant challenges: performance is highly sensitive to the complex engineering of state and reward functions; the resulting policies are often opaque “black boxes” hindering trust; and they demand accurate simulation environments for training, which are often unavailable for complex real-world systems.
@@
Large Language Models Paradigms.
Seeking to address the challenges of both handcrafted logic and the black-box nature of DRL, LLMs have recently introduced a new paradigm for scheduling by leveraging their vast pre-existing knowledge and reasoning capabilities. One direction explores LLMs as end-to-end solvers by fine-tuning them on large-scale datasets of problems paired with their solutions (Abgaryan, Cazenave, and Harutyunyan 2025; Abgaryan, Harutyunyan, and Cazenave 2024). Another approach employs LLMs as automated problem modelers that translate natural language into executable code for traditional solvers (Amarasinghe et al. 2023; AhmadiTeshnizi et al. 2024). A third direction positions the LLM as a hyper-heuristic generator that iteratively reflects upon and improves heuristic rule-code (Huang et al. 2024). Despite their potential, early works reveal critical challenges: the risk of generating infeasible solutions (“hallucinations”), high computational costs, and a primary focus on static problems, leaving their application to complex, dynamic scheduling environments largely unexplored.
3 Preliminaries
Dynamic Flexible Job-Shop Scheduling.
DFJSP is the problem of assigning operations from multiple jobs to a set of machines, subject to real-time stochastic events. It is characterized by two features: flexibility, which allows each operation to be processed on a subset of candidate machines, and dynamism, which refers to the occurrence of real-time stochastic events such as new job arrivals or machine breakdowns. The objective is to generate a valid schedule that minimizes the makespan, defined as the maximum completion time among all jobs (Cao and Yuan 2024). Figure 1 illustrates an example of such a schedule, visualized as a Gantt chart. A valid schedule must satisfy both precedence constraints between operations of the same job and the resource constraints of the machines. A formal mathematical formulation is presented in Appendix A.
Baseline LLM-based Scheduler.
@@ 3 Preliminaries
To evaluate our proposed method, we define a direct baseline, LLM-Direct, which employs the LLM to perform the function of classical dispatching rules. In this approach, a state-aware prompt is constructed at each decision point. The shop-floor state is encoded into the prompt, including both static specifications, such as job structures and machine data, and dynamic information, such as the set of currently available actions and current machine statuses. The action selected by the LLM is then executed in a shop-floor simulator, advancing the system state to the next decision point. Further implementation details for LLM-Direct are presented in Appendix B.
4 Motivational Analysis: The Pitfalls of Direct LLM-based Scheduling
@@ 4.1 The Long-Context Paradox
While replacing state engineering with a single LLM prompt is appealing, it creates an information utility paradox: encoding the high-dimensional shop-floor state into a dense, verbose prompt leads to underutilization of the provided data (Jiang et al. 2024). We observe this is not merely a context-length issue but a fundamental challenge in reasoning over such textual complexity (Liu et al. 2023), as exemplified by our case study with smaller models (Appendix C).
@@ 4.2 Underutilization of Heuristics
A key challenge in LLM-based scheduling is the effective application of PDRs. While LLMs are generally proficient at following declarative instructions, their capacity to faithfully execute complex, state-dependent procedural rules remains unreliable (Mu et al. 2024).
To isolate and measure this capability, we developed PDR-Bench, a diagnostic benchmark where each instance is programmatically constructed to have a single, known optimal heuristic, and employed it in a controlled experiment to evaluate our baseline LLM (see Appendix D for detailed experimental details). The “optimal heuristic” for any given instance in PDR-Bench is known by design, as each problem instance is specifically constructed to ensure that one particular heuristic yields a superior outcome over all others. The results (Figure 3) reveal the LLM’s significant difficulty in effectively utilizing these heuristics (Uzunoglu, Safa, and Şahin 2024). This finding strongly suggests the model’s inability to translate the strategic guidance of the provided rules into its decision-making. Furthermore, when explicitly instructed to use the single optimal heuristic, the model still failed to match the performance of that heuristic applied in isolation, reverting instead to its generalized, pre-trained behaviors.
Figure 3: Analysis of heuristic utilization on the PDR-Bench dataset, with Relative Percent Deviation (RPD) measured against the known optimal heuristic for each instance (0% baseline). The baseline LLM struggles to apply the optimal heuristic even when explicitly prompted, whereas ReflecSched substantially improves performance.

### v2

@@ Abstract
The NP-hard Dynamic Flexible Job-Shop Scheduling (DFJSP) problem involves real-time events and complex routing. While traditional rules are efficient but rigid, deep learning is opaque and requires feature engineering. Large Language Models (LLMs) promise adaptive reasoning without this engineering overhead, yet we find their direct application is suboptimal. Baseline LLMs suffer from three key pitfalls: the long-context paradox, where crucial data is underutilized; an underutilization of expert heuristics; and myopic decision-making. To address this, we propose ReflecSched, a framework that empowers the LLM beyond a direct scheduler by equipping it with a strategic analysis capability. ReflecSched tasks the LLM to analyze heuristic-driven simulations across multiple planning horizons and distill them into a concise, natural-language summary termed “Strategic Experience”. This summary is then integrated into the prompt of a final decision-making module, guiding it to produce non-myopic actions. Experiments demonstrate ReflecSched achieves superior performance, with its best variants attaining an average RPD of 6.04% and rank of 3.18, significantly outperforming strong traditional and learning-based methods. It also statistically and decisively surpasses direct LLM baselines, securing a 71.35% Win Rate while being, on average, 15.1% more token-efficient on Normal-scale problems. Ablation studies attribute this performance to a robust reflection mechanism that leverages high-quality, contrastive experience. This mechanism mitigates key LLM pitfalls like myopic greed, enabling ReflecSched to outperform all evaluated heuristics. Ultimately, the framework’s performance is statistically on par with an oracle-like strategy, showcasing its effectiveness and robustness.
keywords:
Dynamic Flexible Job-Shop Scheduling , Large Language Models , Hierarchical Reflection , Deep Reinforcement Learning , LLM-based Planning
journal: Journal of Manufacturing Systems
@@ 1 Introduction
Dynamic Flexible Job-Shop Scheduling (DFJSP) is a long-standing challenge in operations research and a key component of modern smart manufacturing [li2025learning]. As an NP-hard problem, it involves the complex task of continuously allocating operations to machines in a real-time environment characterized by stochastic events such as new job arrivals or machine breakdowns. The ability to efficiently schedule these operations critically impacts the agility, resilience, and profitability of production systems. For decades, the field has been dominated by methods ranging from classical heuristics to meta-heuristics [cao2023inverse]. While foundational, these approaches often rely on handcrafted rules that can struggle to generalize across diverse and unforeseen dynamic scenarios [ferreira2022effective].
In recent years, deep learning (DL), particularly reinforcement learning (RL), has emerged as a promising paradigm for this problem [xu2025learn]. By training agents on simulated data, DL-based methods can learn sophisticated, state-aware scheduling policies. However, this capability comes at a significant cost. A key limitation lies in the intricate process of state and action space engineering: translating the rich state of a dynamic workshop into a fixed-size numerical vector that a neural network can process is a complex task requiring dual expertise in scheduling and machine learning [zhang2020learning]. Furthermore, the resulting models are often “black boxes” with opaque decision-making logic that is difficult for engineers to interpret or trust in high-stakes environments [ngwu2025reinforcement].
The advent of Large Language Models (LLMs) offers a different approach, potentially circumventing these challenges by leveraging natural language as a more accessible interface. An LLM-based scheduler could, in principle, reason directly over a textual description of the factory state, reducing the need for complex numerical encoding [abgaryan2024llms]. This suggests the possibility of developing and adapting scheduling strategies with greater flexibility.
@@
  • 1.
    A rigorous empirical diagnosis of LLM-based schedulers. We are the first to systematically identify and validate three fundamental failure modes of LLMs in the complex DFJSP domain: the long-context paradox, heuristic underutilization, and myopic greed. This analysis provides crucial insights for the broader field of LLM-based planning.
  • 2.
    A novel LLM-based scheduling paradigm, ReflecSched. We propose a framework that recasts the LLM from a reactive decision-maker into a strategic analyst. By decoupling long-horizon planning from immediate execution via a hierarchical reflection mechanism, ReflecSched is designed to mitigate the identified pitfalls.
@@ 2 Related Work
Heuristic and Metaheuristic Approaches: Heuristic-based methods are a foundational approach for the Dynamic Job-Shop Scheduling Problem (DJSP) due to their computational tractability. Research has advanced from simple dispatching rules to complex metaheuristics designed to handle diverse operational constraints. For instance, some works have applied GRASP algorithms to manage a wide array of dynamic events [baykasouglu2020greedy], while others have used hybrid Particle Swarm Optimization for joint production and transportation scheduling [ren2022joint]. Further research has focused on specific uncertainties, such as variable processing times, using techniques like Artificial Bee Colony algorithms [shahgholi2019heuristic]. A key limitation of these metaheuristics, however, is their reliance on a pre-defined search logic that is inherently less adaptive and struggles to generalize to new, unseen problem scenarios.
@@
Deep Reinforcement Learning Approaches: To address the generalization limitations of heuristics, Deep Reinforcement Learning (DRL) has emerged as an alternative that learns adaptive policies, typically by formulating the DJSP as a Markov Decision Process. Architectural innovations have been central to this area. For example, Pointer Networks have been utilized for end-to-end policy learning that avoids reliance on predefined rules [yang2024deep]. To enhance scalability, a multi-agent perspective has been adopted through decentralized frameworks [liu2023deep]. Other approaches decompose the decision-making process using hierarchical structures [luo2021real]. Nevertheless, DRL presents its own significant challenges: performance is highly sensitive to the complex engineering of state and reward functions; the resulting policies are often opaque “black boxes” hindering trust; and they demand accurate simulation environments for training, which are often unavailable for complex real-world systems.
Large Language Models Paradigms: Seeking to address the challenges of both handcrafted logic and the black-box nature of DRL, LLMs have recently introduced a new paradigm for scheduling by leveraging their vast pre-existing knowledge and reasoning capabilities. One direction explores LLMs as end-to-end solvers by fine-tuning them on large-scale datasets of problems paired with their solutions [abgaryan2025starjob, abgaryan2024llms]. Another approach employs LLMs as automated problem modelers that translate natural language into executable code for traditional solvers [amarasinghe2023ai, ahmaditeshnizi2024optimus]. A third direction positions the LLM as a hyper-heuristic generator that iteratively reflects upon and improves heuristic rule-code [huang2024automatic]. Despite their potential, early works reveal critical challenges: the risk of generating infeasible solutions (“hallucinations”), high computational costs, and a primary focus on static problems, leaving their application to complex, dynamic scheduling environments largely unexplored.
Hierarchical Memory and Reasoning Paradigms: A significant body of work explores hierarchical paradigms to structure and access information beyond the fixed context window. A prominent direction involves creating hierarchical memory systems, such as MemGPT’s OS-inspired information paging [packer2024memgptllmsoperatingsystems] or the semantic-level memory structures in MemoryBank and H-MEM that enable targeted retrieval [zhong2023memorybankenhancinglargelanguage, sun2025hierarchicalmemoryhighefficiencylongterm]. A related approach, seen in HIAGENT, employs hierarchy for task decomposition to manage an agent’s working memory during long-horizon tasks [hu2024hiagenthierarchicalworkingmemory]. Diverging from these methods, which primarily focus on organizing and retrieving an agent’s past, our proposed ReflecSched framework introduces a hierarchical reflection mechanism that functions as a dynamic, planning-time operator. Instead of retrieving historical data, it prospectively explores the future by leveraging multi-level, heuristic-driven simulations to analyze potential decision pathways. The hierarchy in ReflecSched is thus one of planning abstraction distilling numerical simulations into a transient, textual Strategic Experience to guide the next action, rather than creating a persistent memory store.
3 Preliminaries
3.1 Dynamic Flexible Job-Shop Scheduling
DFJSP is the problem of assigning operations from multiple jobs to a set of machines, subject to real-time stochastic events. It is characterized by two features: flexibility, which allows each operation to be processed on a subset of candidate machines, and dynamism, which refers to the occurrence of real-time stochastic events such as new job arrivals or machine breakdowns. The objective is to generate a valid schedule that minimizes the makespan, defined as the maximum completion time among all jobs [cao2024novel]. Figure 1 illustrates an example of such a schedule, visualized as a Gantt chart. A valid schedule must satisfy both precedence constraints between operations of the same job and the resource constraints of the machines.
In this work, we model a rich set of dynamic events to capture the stochastic nature of modern manufacturing environments. These include new job arrivals, machine breakdowns and subsequent repairs, job cancellations, and the introduction of high-priority emergency jobs. This set of events is representative of the primary disruptions studied in the DFJSP literature [baykasoglu2020greedy, ngwu2025reinforcement]. More importantly, the core architecture of ReflecSched is designed to be event-agnostic and generalizable. Its strategic Hierarchical Reflection module is invoked by an event-driven trigger, which responds to any significant state change rather than to a specific event type. Consequently, the framework can be readily extended to handle other stochastic events, such as variable processing times [shahgholi2019heuristic] or rush order arrivals [ren2024dynamic], without requiring modifications to its fundamental logic.
@@ 3.1 Dynamic Flexible Job-Shop Scheduling
3.2 Mathematical Formulation of the DFJSP
The notation used throughout is summarized in Table 1.
Table 1: Notation for the DFJSP Mathematical Formulation
Symbol Description
$\mathcal{J}$ Set of active jobs currently in the system, $\{1,...,n\}$.
$\mathcal{M}$ Set of available machines, $\{1,...,m\}$.
$\mathcal{O}_{i}$ Set of operations for job $J_{i}\in\mathcal{J}$.
$\mathcal{O}_{int}$ Set of interrupted operations awaiting machine repair.
$\mathcal{O}$ Set of all operations to be scheduled.
$M_{ij}$ Set of candidate machines for operation $O_{ij}$.
$n_{i}$ Number of operations for job $J_{i}$.
$p_{ijk}$ Processing time of $O_{ij}$ on machine $k\in M_{ij}$.
$a_{i}$ Arrival time of job $J_{i}$.
$r_{k}(t)$ Earliest time machine $k$ is free.
$M$ A sufficiently large positive number (a “Big-M” constant).
$s_{ij}$ Start time of operation $O_{ij}$.
$c_{ij}$ Completion time of operation $O_{ij}$.
$x_{ijk}$ Binary: 1 if $O_{ij}$ is assigned to machine $k$.
$y_{ij,i^{\prime}j^{\prime}k}$ Binary: 1 if operation $O_{ij}$ precedes $O_{i^{\prime}j^{\prime}}$ when both are assigned to machine $k$.
$C_{max}$ The makespan, defined as $\max_{i\in\mathcal{J}}\{c_{i,n_{i}}\}$.
3.2.1 Objective Function and Constraints
The primary objective is to minimize the makespan ($C_{max}$), which is defined as the maximum completion time among all jobs.
$\min C_{max}$ (1)
subject to:
$C_{max}\geq c_{i,n_{i}}\quad\forall J_{i}\in\mathcal{J}$ (2)
The schedule must satisfy the following constraints:
Processing Time: The completion time of an operation is determined by its start time and the processing time on its assigned machine.
$c_{ij}-s_{ij}=\sum_{k\in M_{ij}}p_{ijk}\cdot x_{ijk}\quad\forall O_{ij}\in\mathcal{O}$ (3)
Assignment: Each operation must be assigned to exactly one of its candidate machines.
$\sum_{k\in M_{ij}}x_{ijk}=1\quad\forall O_{ij}\in\mathcal{O}$ (4)
Job Arrival: An operation cannot start before its corresponding job has arrived.
$s_{ij}\geq a_{i}\quad\forall O_{ij}\in\mathcal{O}$ (5)
Precedence: Within each job, operations must be performed in their specified sequence.
$s_{i,j+1}\geq c_{ij}\quad\forall J_{i}\in\mathcal{J},\forall j\in\{1,\dots,n_{i}-1\}$ (6)
Machine Capacity: For any two distinct operations, if they are assigned to the same machine, they cannot be processed simultaneously. This is enforced by the following disjunctive constraints for any pair of operations $(O_{ij},O_{i^{\prime}j^{\prime}})$ where $i\neq i^{\prime}$ or $j\neq j^{\prime}$, and for each machine $k\in M_{ij}\cap M_{i^{\prime}j^{\prime}}$:
$s_{i^{\prime}j^{\prime}}$ $\geq c_{ij}-M\cdot(3-x_{ijk}-x_{i^{\prime}j^{\prime}k}-y_{ij,i^{\prime}j^{\prime}k})$ (7)
$s_{ij}$ $\geq c_{i^{\prime}j^{\prime}}-M\cdot(2+y_{ij,i^{\prime}j^{\prime}k}-x_{ijk}-x_{i^{\prime}j^{\prime}k})$ (8)
3.2.2 Formalization of Dynamic Events
The dynamic nature of the DFJSP is addressed by updating this formulation in response to stochastic events. Let the state of the system at time $t$ be defined by the tuple $\Sigma_{t}=(\mathcal{J}_{t},\{a_{i}\},\{r_{k}(t)\},\mathcal{O}_{int})$. A dynamic event at $t_{event}$ triggers a state transition from $\Sigma_{t_{event}}$ to $\Sigma_{t_{event}^{+}}$, where $t_{event}^{+}$ denotes the time immediately after the event. The updates for each event type are formalized below [baykasouglu2020greedy].
Job Arrival: At time $t_{event}=t_{arr}$, a new job $J_{new}$ with its own set of operations $\mathcal{O}_{new}$ and arrival time $a_{new}=t_{arr}$ is introduced into the system. The state is updated as follows:
$\mathcal{J}_{t_{arr}^{+}}$ $\leftarrow\mathcal{J}_{t_{arr}}\cup\{J_{new}\}$ (9)
$\mathcal{O}_{t_{arr}^{+}}$ $\leftarrow\mathcal{O}_{t_{arr}}\cup\mathcal{O}_{new}$ (10)
The emergency status of a job is a characteristic handled by the scheduling policy (e.g., via prioritization) rather than being a direct modification to the model’s constraints.
Job Cancellation: At time $t_{event}=t_{cancel}$, a job $J_{cancel}$ is removed from the system. This action removes the job from the set of active jobs $\mathcal{J}$ and all of its corresponding uncompleted operations from the set of operations to be scheduled $\mathcal{O}$. The historical record of any completed operations of $J_{cancel}$ is preserved, but they are no longer relevant for future scheduling decisions.
The state update is formalized as:
$\mathcal{J}_{t_{cancel}^{+}}$ $\leftarrow\mathcal{J}_{t_{cancel}}\setminus\{J_{cancel}\}$ (11)
$\mathcal{O}_{t_{cancel}^{+}}$ $\leftarrow\mathcal{O}_{t_{cancel}}\setminus\mathcal{O}_{cancel}$ (12)
If an operation from $\mathcal{O}_{cancel}$ was in process on a machine $k$ at $t_{cancel}$, it is preempted. The machine is immediately freed, and its availability time $r_{k}(t_{cancel}^{+})$ is updated to $t_{cancel}$.
Machine Breakdown: At time $t_{event}=t_{break}$, a machine $k\in\mathcal{M}$ fails and becomes unavailable. The immediate consequence is that the machine’s earliest availability time is postponed until its projected repair time, $t_{repair}$. This is formalized as:
$r_{k}(t_{break}^{+})=t_{repair}$ (13)
This update affects all subsequent scheduling decisions, as no new operations can be assigned to machine $k$ until $t_{repair}$.
If an operation $O_{ij}$ was being processed on machine $k$ at this time (i.e., $s_{ij}<t_{break}<c_{ij}$), it is immediately interrupted. The operation is formally moved to a set of interrupted operations, $\mathcal{O}_{int}$, to signify its special status. It retains its assignment to machine $k$ and awaits the machine’s repair to resume processing.
Machine Repair: At time $t_{event}=t_{repair}$, machine $k$ becomes operational again. Any operation $O_{ij}\in\mathcal{O}_{int}$ that was interrupted on this machine can now resume. Upon repair, $O_{ij}$ is removed from the set $\mathcal{O}_{int}$.
The resumption of its processing leads to a new, postponed completion time, $c^{\prime}_{ij}$. While the intrinsic processing time $p_{ijk}$ remains unchanged, the new completion time is calculated based on the remaining work:
$c^{\prime}_{ij}=t_{repair}+p_{ijk}-(t_{break}-s_{ij})$ (14)
where $t_{break}-s_{ij}$ represents the processing work completed before the interruption. This new completion time $c^{\prime}_{ij}$ subsequently serves as the basis for re-evaluating the precedence constraints for all downstream operations of job $J_{i}$. The machine $k$ is also now available for scheduling other operations from time $t_{repair}$ onwards.
3.3 Baseline LLM-based Scheduler
@@
To evaluate our proposed method, we define a direct baseline, LLM-Direct, which employs the LLM to perform the function of classical dispatching rules. In this approach, a state-aware prompt is constructed at each decision point. The shop-floor state is encoded into the prompt, including both static specifications, such as job structures and machine data, and dynamic information, such as the set of currently available actions and current machine statuses. The action selected by the LLM is then executed in a shop-floor simulator, advancing the system state to the next decision point. Further implementation details for LLM-Direct are presented in A.
4 Motivational Analysis
@@ 4.1 The Long-Context Paradox
While replacing state engineering with a single LLM prompt is appealing, it creates an information utility paradox: encoding the high-dimensional shop-floor state into a dense, verbose prompt leads to underutilization of the provided data [jiang2024longllmlinguaacceleratingenhancingllms]. We observe this is not merely a context-length issue but a fundamental challenge in reasoning over such textual complexity [liu2023lostmiddlelanguagemodels], as exemplified by our case study with smaller models in B.
@@ 4.2 Underutilization of Heuristics
A key challenge in LLM-based scheduling is the effective application of PDRs. While LLMs are generally proficient at following declarative instructions, their capacity to faithfully execute complex, state-dependent procedural rules remains unreliable [mu2024llmsfollowsimplerules].
To isolate and measure this capability, we developed PDR-Bench, a diagnostic benchmark where each instance is programmatically constructed to have a single, known optimal heuristic, and employed it in a controlled experiment to evaluate our baseline LLM (see C for detailed experimental details). The “optimal heuristic” for any given instance in PDR-Bench is known by design, as each problem instance is specifically constructed to ensure that one particular heuristic yields a superior outcome over all others. The results (Figure 3) reveal the LLM’s significant difficulty in effectively utilizing these heuristics [uzunoglu2024paradiseevaluatingimplicitplanning]. This finding strongly suggests the model’s inability to translate the strategic guidance of the provided rules into its decision-making. Furthermore, when explicitly instructed to use the single optimal heuristic, the model still failed to match the performance of that heuristic applied in isolation, reverting instead to its generalized, pre-trained behaviors.
Figure 3: Analysis of heuristic utilization on the PDR-Bench dataset, with RPD measured against the known optimal heuristic for each instance (0% baseline). The baseline LLM struggles to apply the optimal heuristic even when explicitly prompted, whereas ReflecSched substantially improves performance.